OpenStudy (anonymous):

Let f(t)=4 t^2 + 4/t^2. Find the slope of the curve at t = -3.

OpenStudy (radar):

f(t)'=8t -(8)/t^3 for t=-3 -24-8/27 -640/27=-20

OpenStudy (radar):

Ooop it is not -20 -640/27=-23.7037

OpenStudy (anonymous):

yeah I got df/dt=8t-2/t^2 so t=-3 => 8*(-3)-2/(-3)^2 -24-2/9 ! but i dont trust my work

OpenStudy (radar):

the second term when you differentiated;\[4 \over t ^{2}\]is\[4 t ^{-2}\]becomes \[-8 t ^{-3}\]

OpenStudy (radar):

or\[-8\over t ^{3}\]

OpenStudy (anonymous):

but I would still have to plug in -3 for t

OpenStudy (radar):

Yes indeed.

OpenStudy (anonymous):

wait was I on the right track to getting the answer?

OpenStudy (radar):

that would be -8/27 for that term and 24 for the first term I see what you mean it can't be negative with 24 being positive!! lol, I better take a nap.

OpenStudy (anonymous):

but -640/27=-23.7037 is the answer

OpenStudy (anonymous):

right?

OpenStudy (radar):

Oh yes, thats right as t=-3 so it would be -24 and a negative 8/27 that means we have to combine the -24 and the -8/27 giving -648/27 - 8/27=-656/27= -24.29

OpenStudy (radar):

Looking at your next to last post. How did you get -640/27? Wouldn't -24 X 27 =-648 ??

OpenStudy (radar):

It would appear so, if I have kept track of the signs.

OpenStudy (radar):

But maybe I messed up let me think about this?

OpenStudy (anonymous):

lol take your time

OpenStudy (radar):

I think I did mess up on that 2nd term when I plugged in the minus 3\[-8\over t ^{3}\]\[-8\over (-3)^{3}\]\[-8\over -27\] or POSITIVE 8/27 lol there it is.

OpenStudy (radar):

So we ADD the 8/27 to the negative 648/27 giving -640/27 So sorry paunic88, I did stray off the path of keeping track of signs.

OpenStudy (anonymous):

ok so the final anwser is -640/27 right? (ill compute it later)

OpenStudy (radar):

-3 X -3 X -3 = -27 lol

OpenStudy (radar):

Yes.

OpenStudy (radar):

Good luck with your studies,and don't err with signs lol

OpenStudy (anonymous):

ok thank you very much for the help

OpenStudy (radar):

ur welcome.